### v1

# Benchmarking Filtered Approximate Nearest Neighbor Search Algorithms on Transformer-based Embedding Vectors

Abstract.
Advances in embedding models for text, image, audio, and video drive progress across multiple domains, including retrieval-augmented generation, recommendation systems, vehicle/person re-identification, and face recognition. Many applications in these domains require an efficient method to retrieve items that are close to a given query in the embedding space while satisfying a filter condition based on the item’s attributes, a problem known as Filtered Approximate Nearest Neighbor Search (FANNS). In this work, we present a comprehensive survey and taxonomy of FANNS methods and analyze how they are benchmarked in the literature. By doing so, we identify a key challenge in the current FANNS landscape: the lack of diverse and realistic datasets, particularly ones derived from the latest transformer-based text embedding models. To address this, we introduce a novel dataset consisting of embedding vectors for the abstracts of over 2.7 million research articles from the arXiv repository, accompanied by 11 real-world attributes such as authors and categories. We benchmark a wide range of FANNS methods on our novel dataset and find that each method has distinct strengths and limitations; no single approach performs best across all scenarios. ACORN, for example, supports various filter types and performs reliably across dataset scales but is often outperformed by more specialized methods. SeRF shows excellent performance for range filtering on ordered attributes but cannot handle categorical attributes. Filtered-DiskANN and UNG excel on the medium-scale dataset but fail on the large-scale dataset, highlighting the challenge posed by transformer-based embeddings, which are often more than an order of magnitude larger than earlier embeddings. We conclude that no universally best method exists.
Code:
https://github.com/spcl/fanns-benchmark
Datasets:
https://huggingface.co/datasets/SPCL/arxiv-for-fanns-small https://huggingface.co/datasets/SPCL/arxiv-for-fanns-medium https://huggingface.co/datasets/SPCL/arxiv-for-fanns-large
1. Introduction
Advances in embedding models for text (Lee et al., 2024; Lei et al., 2025; Zhang et al., 2024), image (Szegedy et al., 2015; Liu et al., 2024), video (Sivic and Zisserman, 2003a), audio (Cramer et al., 2019; Aytar et al., 2016), and other modalities (Lerer et al., 2019) have significantly enhanced semantic search, where items are mapped to a high-dimensional vector space, and similarity is measured by the distance between embedding vectors. Efficient similarity search algorithms are essential to navigate this space. Given the scale of modern datasets, exact nearest neighbor search (ENNS) algorithms are too slow, necessitating the use of approximate nearest neighbor search (ANNS) methods (Indyk and Motwani, 1998; Malkov et al., 2014; Malkov and Yashunin, 2018; Jayaram Subramanya et al., 2019; Sivic and Zisserman, 2003b; Jegou et al., 2010; Gong et al., 2020; Babenko and Lempitsky, 2014a; Guo et al., 2016; Ge et al., 2013).
Many real-world applications, including large language models (LLMs) with retrieval-augmented generation (Gao et al., 2023; Xu et al., 2021; Shapkin et al., 2023), recommendation systems (Suchal and Návrat, 2010; Yang et al., 2020b; Koenigstein et al., 2012), vehicle and person re-identification (Liu et al., 2016a; Zuo and Deng, 2023), face and voice recognition (Grother et al., 2019; Yang et al., 2020b; Richardson et al., 2015), and e-commerce (Li et al., 2018; Wei et al., 2020), require retrieving only items that satisfy specific filtering conditions on item attributes, e.g., access rights for a document, a video’s timestamp, or a product’s price. These requirements have driven the development of filtered approximate nearest neighbor search (FANNS). Both academia and industry have recognized its growing importance, leading to numerous research publications (Wang et al., 2024; Mohoney et al., 2023; Gollapudi et al., 2023; Gupta et al., 2023; Zuo and Deng, 2023; Patel et al., 2024; Engels et al., 2024; Zuo et al., 2024; Xu et al., 2024; Liang et al., 2024) and adoption in database solutions such as Pinecone (Pinecone, 2025), Vespa (Vespa, 2025), and Vectara (Vectara, 2025). Evaluating FANNS methods requires vector datasets that reflect real-world workloads. Yet, despite the rapid rise of transformer-based text embeddings, no FANNS benchmarks with such vectors are known to us.
Our contributions are threefold. First, to clarify the requirements for effective FANNS benchmarks, we survey the current FANNS landscape and propose a comprehensive taxonomy of existing methods. We classify approaches along three key dimensions: filtering approach, indexing technique, and supported filter types. Second, we present a novel dataset featuring vectors generated by embedding paper abstracts from the arXiv dataset (Cornell University et al., 2025) using the transformer-based model stella_en_400M_v5 (NovaSearch, 2025; Zhang et al., 2024). Each item in our dataset has 111111 real-world attributes. Third, we benchmark a diverse set of FANNS methods on this dataset. Our evaluation reveals that each method involves distinct trade-offs, and no single approach dominates across all conditions. For example, ACORN supports multiple filter types and scales well, but is often outperformed by more specialized methods. Filtered-DiskANN and UNG perform strongly at medium scale but fail on our largest dataset. SeRF performs consistently well but supports only range filters on ordered attributes. These results demonstrate that FANNS remains a nuanced design space, and that selecting the right method depends on the specific workload and filtering requirements.
2. Background
2.1. FANNS Problem Statement
In FANNS, we are given $n$ items, $m\geq 1$ attributes, and $p$ queries. Each item $I_{i}=(v_{i},a_{i,1},...,a_{i,m})$ for $i\in\{1,\dots,n\}$ consists of a $d$-dimensional embedding vector $v_{i}$ and a value for each of the $m$ attributes ($a_{i,1}$ to $a_{i,m}$). An item can, e.g., represent a research paper with $v_{j}$ being an embedding of its abstract and attributes corresponding to venue, year of publication, and authors. A query $Q_{j}=(q_{j},k_{j},f_{j})$ for $j\in\{1,\dots,p\}$ consists of a $d$-dimensional query vector $q_{j}$, the number of items to return $k_{j}$, and a filter function $f_{j}$ that determines whether an item matches the filter. For a query, $q_{j}$ can, e.g., be the embedding of a search term with $f_{j}$ only returning true for papers submitted to VLDB. We use $o_{j}\leq m$ to denote the number of attributes the filter $f_{j}$ depends on. Table 1 summarizes all parameters.
Table 1. 2.1) Parameters used in FANNS.
Parameter Description
$n$ Number of items
$m$ Number of attributes
$p$ Number of queries
$d$ Dimensionality of embedding and query vectors
$v_{i}$ Embedding vector of the $i$-th item
$a_{i,l}$ $l$-th attribute of the $i$-th item
$q_{j}$ Query vector of the $j$-th query
$k_{j}$ Number of items to return for the $j$-th query
$f_{j}$ Filter expression of the $j$-th query
$o_{j}$ Number of attributes considered in $f_{j}$
To answer a query $Q_{j}$, we approximately retrieve the $k_{j}$ items that satisfy the filter $f_{j}$ and have embedding vectors closest to the query vector $q_{j}$ according to a given distance function (e.g., Euclidean distance, cosine similarity, Manhattan distance, or Hamming distance). Most FANNS algorithms return only items that match the filter; however, some approximate the filtering step, occasionally retrieving items that do not satisfy it. We refer to this relaxed version of the FANNS problem as the approximately filtered approximate nearest neighbor search (AFANNS) problem.
2.2. Example Use-Case of FANNS
To illustrate the practical use of FANNS, we introduce an example application depicted in Figure 1. We reference components and actions in Figure 1 using letters [Uncaptioned image] and numbers [Uncaptioned image], respectively. Consider a semantic search engine for research papers; a system that retrieves the most relevant papers for a given query, even when different terminology is used to describe the same concept. To provide fine-grained control over search results, we incorporate filters for venue, publication year, and authors.
Index Construction: To build such a system, we transform each paper in our database [Uncaptioned image] into an item, as described in Section 2.1. The paper’s venue, publication year, and authors define [Uncaptioned image] the item’s attributes. The paper’s abstract is processed [Uncaptioned image] by a text embedding model [Uncaptioned image], such as NV-Embed (Lee et al., 2024; NVIDIA, 2025), LENS (Lei et al., 2025; Lei, 2025), Stella (Zhang et al., 2024; NovaSearch, 2025), GTE (Li et al., 2023; Alibaba-NLP, 2025), or BGE (Chen et al., 2024; of Artificial Intelligence, 2025), to generate the item’s embedding vector. To efficiently answer FANNS queries for these items, we insert [Uncaptioned image] them into a FANNS index [Uncaptioned image].
Query execution: When a user enters [Uncaptioned image] a search query, we transform it into a FANNS query, as described in Section 2.1. The filter $f_{j}$ and the number of requested results $k_{j}$ are set [Uncaptioned image] based on the user’s input. The query text is processed [Uncaptioned image] by the same text embedding model [Uncaptioned image] used during index construction to generate the query vector. The query is submitted [Uncaptioned image] to a FANNS algorithm [Uncaptioned image], which utilizes [Uncaptioned image] the FANNS index [Uncaptioned image] to retrieve the IDs of the most relevant papers. These IDs are forwarded [Uncaptioned image] to a retriever [Uncaptioned image], which fetches [Uncaptioned image] the corresponding papers from the database [Uncaptioned image] and returns [Uncaptioned image] them to the user.
Figure 1. 2.2) Example use case of FANNS.
2.3. Attribute Types
We provide a list of attribute types supported by FANNS algorithms.
  • Unordered attributes: These attributes contain a single value from a domain with no total order. An example is the venue of a published paper, as each paper appears in a single venue, and no total order exists between venues.
  • Ordered attributes: These attributes contain a single value from a totally ordered domain. An example is the publication year of a paper, as each paper has a single publication year, and years follow a total order.
  • Set attributes: These attributes contain a set of values from an unordered domain. For instance, the authors of a research paper can be represented as a set attribute.
2.4. Filter Types
We categorize filters that apply to a single attribute into three types. Table 2 shows the compatibility of filters types with attribute types.
  • Exact match (EM) filter: Matches only if an item’s attribute value is equal to the queried value.
  • Range (R) filter: Used on an ordered attribute and matches only if an item’s attribute value is within the queried range.
  • Exact match in set (EMIS) filter: Applies to a set attribute and matches only if the item’s set attribute contains the queried value.
FANNS algorithms that apply filtering to multiple attributes of the same type implement multiple exact match (MEM), multiple range (MR), and multiple exact match in set (MEMIS) filters. We define a combined (C) filter as a filter capable of combining the three basic single-attribute filters using arbitrary logical operators.
Table 2. 2.4) Which filter applies to which attribute.
Unordered attribute Ordered attribute Set attribute
Exact match filter
Range filter
Exact match in set filter
2.5. Approximate Nearest Neighbor Search
Since many FANNS algorithms build on ANNS methods without filtering, we first introduce the most relevant ANNS algorithms. These can be categorized into tree-based, hash-based, graph-based, and quantization-based methods. For a comprehensive discussion of ANNS, we refer to the surveys by Han et al. (Han et al., 2023) and Echihabi et al. (Echihabi et al., 2021). Experimental evaluations (Aumüller et al., 2020; Aumueller et al., 2025) indicate that no single ANNS algorithm consistently outperforms others; rather, performance depends heavily on dataset characteristics.
2.5.1. Tree-based methods
Tree-based methods partition the vector space into multiple regions by constructing a search tree, where nodes represent either regions or boundaries between them. Lower tree levels typically correspond to finer partitioning of the vector space. Tree-based methods, such as k-d trees (Bentley, 1975), R-trees (Guttman, 1984), ball trees (Omohundro, 1989), cover trees (Beygelzimer et al., 2006), RP trees (Dasgupta and Freund, 2008), M-trees (Ciaccia et al., 1997), K-means trees (Tavallali et al., 2021), and best bin first (Beis and Lowe, 1997), along with other variants (Houle and Nett, 2014; Lu et al., 2020; Muja and Lowe, 2014; Silpa-Anan and Hartley, 2008), are more efficient for lower-dimensional vectors, as they often suffer from the curse of dimensionality (Marimont and Shapiro, 1979).
2.5.2. Hash-based methods
Hash-based methods use a set of hash functions to map vectors from a continuous $d$-dimensional space into discrete hash buckets. These functions are designed so that similar vectors are mapped into the same bucket with high probability. When querying a hash-based index, the query vector is hashed, and the vectors in the corresponding bucket are compared to the query vector. Hash-based methods include locality-sensitive hashing (LSH) (Indyk and Motwani, 1998; Datar et al., 2004), spectral hashing (Weiss et al., 2008), iDEC (Gong et al., 2020), deep hashing (Liu et al., 2016b), mmLSH (Jafari et al., 2020), PM-LSH (Zheng et al., 2020), R2LSH (Lu and Kudo, 2020), EI-LSH (Liu et al., 2021), and others (Andoni et al., 2015; Andoni and Razenshteyn, 2015; Gionis et al., 1999; Li et al., 2020; Lv et al., 2017; Park et al., 2015; Sundaram et al., 2013; Gan et al., 2012; Huang et al., 2015).
2.5.3. Graph-based methods
Graph-based methods represent vectors as vertices in a graph. Two vertices are connected if their corresponding vectors are close to each other, forming a neighborhood graph (Wang and Li, 2012; Aoyama et al., 2011). For simplicity, we define the length of an edge as the distance between its endpoints’ vectors. To find the nearest neighbors of a query vector, the graph is traversed starting from one or multiple entry points, following the direction of the smallest distance to the query vector. A drawback of neighborhood graphs is that for large datasets, many edges must be traversed, leading to high query times. Navigable small world (NSW) graphs (Malkov et al., 2014) mitigate this issue by introducing long edges that significantly reduce the number of steps needed for traversal. Hierarchical navigable small world (HNSW) graphs (Malkov and Yashunin, 2018) further improve the performance of NSW by introducing a hierarchical structure, where higher layers contain longer edges and lower layers contain shorter ones. The HNSW graph is traversed from the highest to the lowest level. Additional graph-based ANNS methods include DiskANN (Jayaram Subramanya et al., 2019), FreshDiskANN (Singh et al., 2021), NSG (Fu et al., 2017), GRNG (Foster et al., 2025), and others (Zhao et al., 2020; Fu et al., 2021; Harwood and Drummond, 2016; Iwasaki, 2016; Fu et al., 2019).
2.5.4. Quantization-based methods
[…]
3. Survey and Taxonomy of FANNS
Table 3. 3) Overview of FANNS methods. Does not mention a filtering step during query execution. *We have not been able to find the source code of this work. [Uncaptioned image] The index is constructed for a single ordered attribute, in-filtering is proposed to handle secondary attributes. [Uncaptioned image] Any EMIS filter can be used as an EM filter by using a set attribute with only a single value.
Filters Open
Method Year Problem Approach Technique(s)
EM
R
EMIS
MEM
MR
MEMIS
C
source
Rii (Matsui et al., 2018) 2018 FANNS pre-/in-filtering
quantization
(Matsui, 2025)
MA-NSW (Xu et al., 2020) 2019 FANNS hybrid index
graph
PASE (Yang et al., 2020b) 2020 FANNS post-filtering
quantization / graph
(Alipay, 2025)
AnalyticDB-V (Wei et al., 2020) 2020 FANNS pre-/post-filtering
quantization
Milvus (Wang et al., 2021d) 2021 FANNS pre-/post-filtering
quantization / graph
(Project, 2025)
NHQ (Wang et al., 2024, 2022) 2022 AFANNS hybrid index
graph
(Fu, 2025)
HQANN (Yang et al., 2020c) 2022 AFANNS hybrid index
graph
AIRSHIP (Zhao et al., 2022) 2022 FANNS in-filtering
graph
HQI (Mohoney et al., 2023) 2023 FANNS hybrid index
tree + quantization
VBASE (Zhang et al., 2023) 2023 FANNS post-filtering
any
(Microsoft, 2025b)
FDANN (Gollapudi et al., 2023) 2023 FANNS hybrid index
graph
[Uncaptioned image] (Microsoft, 2025a)
CAPS (Gupta et al., 2023) 2023 FANNS hybrid index
quantization + tree
(Gupta, 2025)
ARKGraph (Zuo and Deng, 2023) 2023 FANNS hybrid index
graph
(at Rutgers University, 2025a)
ACORN (Patel et al., 2024) 2024 FANNS in-filtering
graph
(Guestrin Lab at Stanford University, 2025)
$\beta$-WST (Engels et al., 2024) 2024 FANNS hybrid index
tree + graph
(Engels, 2025)
SeRF (Zuo et al., 2024) 2024 FANNS hybrid index
graph
[Uncaptioned image] (at Rutgers University, 2025b)
iRangeGraph (Xu et al., 2024) 2024 FANNS hybrid index
tree + graph
[Uncaptioned image] [Uncaptioned image] (Xu, 2025)
UNIFY (Liang et al., 2024) 2024 FANNS hybrid index
graph
(DBGroup, 2025)
UNG (Cai et al., 2024) 2024 FANNS hybrid index
graph
[Uncaptioned image] [Uncaptioned image] (Cai, 2025)
3.1. Filtering Approaches
Two common approaches to solving the FANNS problem are pre-filtering and post-filtering. In pre-filtering, an attribute-only index such as a B-tree (Bayer and McCreight, 1970), B+-tree (Abel, 1984), or qd-tree (Yang et al., 2020a) identifies all items matching the filter. The k nearest neighbors (KNN) of these items are then determined by computing the distance to the query vector or approximating this distance using quantized embedding vectors precomputed during index construction. In post-filtering, a nearest neighbor search is performed on an unfiltered ANNS index, followed by filtering out non-matching items. This can be done by retrieving $k^{\prime}>k$ items in the hope that enough pass the filter or by iteratively retrieving more items until $k$ matches are found. Many graph-based ANNS indices support a third approach, in-filtering, where attributes are ignored during index construction, and only vertices satisfying the filter are considered during query execution. A fourth approach involves a hybrid index that integrates embedding vectors and attributes in a single index.
The efficiency of different approaches to solving the FANNS problem depends on the dataset and filtering condition (see Figure 2). We define the selectivity of a filter $f$ on a dataset $D=\{I_{i}\mid i\in\{1,\dots,n\}\}$ as in previous work [^1] Note that some works (Wei et al., 2020; Wang et al., 2021d; Xu et al., 2024) use the inverse definition of selectivity. (Mohoney et al., 2023; Zhang et al., 2023; Patel et al., 2024; Liang et al., 2024):
$\text{selectivity}=\frac{|\{I_{i}\mid I_{i}\in D\land f(I_{i})=\texttt{true}\}|}{|D|}.$
In other words, selectivity is the fraction of database items that match the filter. Pre-filtering is most efficient when selectivity is low, meaning only a few items remain for the ENNS. Post-filtering and in-filtering are most efficient when selectivity is high, i.e., most items pass the filter. If selectivity is too low, ANNS may not return enough valid candidates for post-filtering, and graph traversal with in-filtering may fail due to a sparsely connected or disconnected graph. A hybrid index is most efficient when selectivity is moderate, making none of the three previous approaches optimal.
3.2. Classification of FANNS Methods
We classify FANNS methods along three dimensions:
  • Approach: Each FANNS method follows one or a combination of the following approaches: pre-filtering, post-filtering, in-filtering, or hybrid indexing (see Section 3.1).
  • Technique: FANNS methods build upon the same four fundamental techniques as ANNS methods: tree-based, hash-based, graph-based, or quantization-based (see Section 2.5).
  • Filter Type: Different FANNS methods support various filter types (see Section 2.4 for details).
Table 3 provides an overview of existing FANNS methods classified along these three dimensions.
Figure 2. 3.1) The optimal approach for solving FANNS depends on the filter’s selectivity within a given dataset.
3.3. Explanation of FANNS Methods
3.3.1. Rii
The Rii (Matsui et al., 2018) index is based on IVF-PQ. It takes a bitmap of matching items as input. If the selectivity is low, it compares the query vector directly with the efficiently accessible, contiguously stored PQ codes (pre-filtering). If the selectivity is high, it first performs the IVF step of IVF-PQ to identify the closest clusters. Within those clusters, a PQ code is compared to the query vector only if the corresponding item matches the filter (in-filtering).
3.3.2. MA-NSW
MA-NSW (Xu et al., 2020) is a graph-based FANNS index that supports both EM and MEM filters. It constructs multiple NSW-like, graph-based ANNS indices, one for each possible combination of attribute values. During query execution, MA-NSW queries the ANNS index that matches the query’s filter expression to approximately retrieve the KNN that satisfy the filter conditions.
3.3.3. PASE
PASE (Yang et al., 2020b) integrates ANNS into the PostgreSQL (Group, 2025) database by implementing the IVF and HNSW ANNS indices. It employs an iterative post-filtering approach, in which it repeatedly retrieves items from the ANNS index and filters them until $k$ matching items are found.
3.3.4. AnalyticDB-V
Alibaba’s AnalyticDB-V (Wei et al., 2020) extends the AnalyticDB (Zhan et al., 2019) SQL database with FANNS capabilities. It introduces its own ANNS index, called Voronoi graph product quantization (VGPQ), which is similar to IVF-PQ. For FANNS implementation, it supports both pre-filtering and post-filtering.
3.3.5. Milvus
Milvus (Wang et al., 2021d) extends the FAISS ANNS library (Douze et al., 2024; Research, 2025) by incorporating range filtering for one or multiple attributes. It supports both pre- and post-filtering, as well as a partition-based filtering scheme that partitions the data based on the most frequently filtered attribute and builds a separate ANNS index for each partition.
3.3.6. NHQ
NHQ (Wang et al., 2024) proposes a method to extend any proximity graph-based ANNS index into a hybrid FANNS index. Instead of constructing the ANNS index solely based on embedding distance, it introduces a fusion distance that combines embedding distance with attribute dissimilarity. While this approach is applicable to any proximity graph-based ANNS index, NHQ also presents construction schemes for two novel proximity graphs. Since its query execution scheme operates using the fusion distance without an explicit filtering step, NHQ may return items that do not match the filter. Thus, it addresses the AFANNS problem rather than the FANNS problem.
3.3.7. HQANN
Like NHQ (Wang et al., 2024), HQANN (Yang et al., 2020c) employs a fusion distance to transform any proximity graph-based ANNS index into a hybrid FANNS index. The key difference between HQANN and NHQ is that HQANN prioritizes attribute dissimilarity in the fusion distance, whereas NHQ emphasizes embedding distance as the dominant factor.
3.3.8. AIRSHIP
AIRSHIP (Zhao et al., 2022) (see Figure 3) is a graph-based FANNS index capable of handling arbitrary filter functions. It relies on a graph-based ANNS index constructed without considering attributes but adapts query execution to FANNS through in-filtering and a series of optimizations. During query execution, AIRSHIP traverses the entire graph but includes only vertices that satisfy the filter in the result set. Additionally, it selects a starting point within a cluster of items guaranteed to match the filter and traverses the graph concurrently in multiple directions.
Figure 3. 3.3.8) Visualization of the AIRSHIP index.
3.3.9. HQI
Apple’s HQI (Mohoney et al., 2023) (see Figure 4) is a hybrid, workload-aware index designed for batch processing of FANNS queries. To integrate attribute filtering with nearest neighbor search, HQI first transforms embedding vectors into attributes by applying k-means clustering and storing each item’s cluster ID as an attribute. It then employs a balanced qd-tree (Yang et al., 2020a), an attribute-only index, to partition the items. Each non-leaf node in the qd-tree contains a predicate over the item attributes (including the cluster ID), with its two child nodes holding items that do or do not satisfy the predicate. The tree’s leaves form a non-overlapping partition of all items. During query execution, HQI identifies the query vector’s $w$ closest cluster IDs and prunes all branches of the tree that do not contain any of these $w$ cluster IDs or whose attributes do not match the filter. An IVF index is used within each leaf.
Figure 4. 3.3.9) Visualization of the HQI index.
3.3.10. VBASE
VBASE (Zhang et al., 2023) extends PostgreSQL (Group, 2025) by integrating ANNS into the database system. Its key insight is that most ANNS query execution methods follow a two-stage process: an initial stage where traversal moves from a starting point in the direction of the query vector, and a second stage where the search gradually moves away from the query vector to identify its KNN. Rather than employing a naïve pre-filtering approach, where the ANNS index is queried with a fixed $k^{\prime}\geq k$ and non-matching items are filtered out, VBASE refines query execution by running the ANNS query only until the traversal starts moving away from the query vector. From that point onward, it iteratively retrieves and filters additional items until the required $k$ matches are found.
3.3.11. FDANN
Microsoft’s Filtered-DiskANN (FDANN) (Gollapudi et al., 2023) (see Figure 5) implements a graph-based index with EMIS filtering. In addition to supporting an EMIS filter with a single query-label, it also allows filtering with multiple query-labels, where at least one of them must be present in the item’s set attribute. Filtered-DiskANN introduces two methods for constructing FANNS indices based on Vamana graphs (Jayaram Subramanya et al., 2019): FilteredVamana and StitchedVamana. During query execution, it traverses the graph starting from a set of entry points that are guaranteed to contain the query-label. Throughout its NSW-like graph traversal, it considers only those vertices that satisfy the filter criteria.
Figure 5. 3.3.11) The Filtered-DiskANN index visualized.
3.3.12. CAPS
In contrast to most FANNS methods, CAPS (Gupta et al., 2023) (see Figure 6) introduces a quantization- and tree-based index that is easier to parallelize than the more common graph-based indices. The CAPS index consists of two levels. The first level organizes items into clusters, for example, using an IVF (Sivic and Zisserman, 2003b). This clustering can be based solely on embedding distance or on a combination of embedding distance and attribute dissimilarity. At the second level, CAPS constructs an attribute frequency tree (AFT), a tree-based attribute-only index. Each level of the AFT splits items into two buckets: one containing items with the most common attribute value (a leaf node) and another with the remaining items (a non-leaf node), unless the maximum depth is reached.
Figure 6. 3.3.12) Visualization of the CAPS index.
During query execution, CAPS first selects the $w$ clusters closest to the query vector (first level). Within these $w$ clusters, it uses AFTs to identify items matching the filter and then performs ENNS on the filtered results (second level).
3.3.13. ARKGraph
[…]
Figure 7. 3.3.13) Visualization of the adjacency lists of item 3 in ARKGraph using the three different approaches.
3.3.14. ACORN
ACORN (Patel et al., 2024) (see Figure 8) supports arbitrary filter types and filtering across multiple attributes. The ACORN index is a denser variant of the HNSW ANNS index (Malkov and Yashunin, 2018), constructed without considering attributes. During query execution, ACORN traverses only the subgraph induced by vertices that match the filter. This approach differs from AIRSHIP (Zhao et al., 2022), which traverses the entire graph but includes only matching vertices in the result set. The ACORN index is designed such that its subgraphs approximate an HNSW index (Malkov and Yashunin, 2018), ensuring efficient nearest neighbor search while maintaining filtering constraints.
Figure 8. 3.3.14) Visualization of the ACORN index.
3.3.15. $\beta$-WST
$\beta$-WST (Engels et al., 2024) (see Figure 9) is a FANNS index designed for range filtering on a single attribute. A $\beta$-WST index is structured as a segment tree (De Berg, 2000) with a uniform branching factor of $\beta$ (see Figure 9). At layer $l$, the attribute value range is divided into $\beta^{l}$ segments, with a graph-based ANNS index constructed for each segment. The authors propose multiple query strategies for $\beta$-WST. The default strategy selects a minimal set of tree nodes that fully cover the query range, queries the corresponding ANNS indices, and merges the results. Additional strategies include OptimizedPostfiltering, TreeSplit, and SuperPostfiltering.
Figure 9. 3.3.15) Visualization of the $\beta$-WST for $\beta=3$.
3.3.16. SeRF
SeRF (Zuo et al., 2024) introduces the segment graph, a data structure designed to efficiently answer range-filtered FANNS queries with half-bounded query ranges (i.e., where the query range has an upper limit but no lower limit). The segment graph losslessly compresses $n$ HNSW indices (Malkov and Yashunin, 2018), corresponding to the $n$ possible half-bounded query ranges, into a single index, which eliminates the need to explicitly construct all $n$ HNSW indices. This is achieved by incrementally inserting items into an HNSW index in attribute-value order. Each edge in the resulting graph stores the attribute value at which it was added and, if pruned during construction, the attribute value at which it was removed. As a result, the HNSW graph contains edges that are valid only within specific attribute value intervals. When executing a query on the segment graph with a half-bounded query range, only edges whose validity interval contains the upper limit of the query range are considered. To generalize this approach for arbitrary query ranges (i.e., with both upper and lower limits), SeRF introduces the 2D segment graph (see Figure 10). This structure compresses $n$ segment graphs, one for each possible lower limit, into a single graph with $n$ vertices, where edges store two validity intervals: one for the lower and one for the upper limit of the query range. Additionally, SeRF proposes an in-filtering approach to extend its method to queries involving two or more attributes.
Figure 10. 3.3.16) SeRF’s 2D segment graph visualized.
3.3.17. iRangeGraph
iRangeGraph (Xu et al., 2024) (see Figure 11) is a graph-based FANNS index designed for range filtering. It is primarily built around a single ordered attribute but also supports in-filtering or post-filtering for secondary attributes of arbitrary types. iRangeGraph constructs a set of graph-based ANNS indices, referred to as elemental graphs. These elemental graphs are organized in a segment tree (De Berg, 2000) with $O(\log n)$ layers. At layer $l$, the entire attribute-value range is divided into $2^{l}$ segments, and a graph-based ANNS index is built for each segment, ensuring that each item appears once in each of the $\log n$ layers. Query execution employs an efficient on-the-fly method to dynamically merge a subset of elemental graphs into a single graph-based ANNS index containing only items within the query range. While its index construction is similar to $\beta$-WST (Engels et al., 2024), iRangeGraph differs significantly in query execution: instead of querying multiple graph-based indices for different subsets of items and merging the results, it constructs a single graph-based ANNS index dynamically at query time.
Figure 11. 3.3.17) Visualization of the iRangeGraph index.
3.3.18. UNIFY
UNIFY (Liang et al., 2024) is a graph-based method that supports range filtering. It introduces two new types of proximity graph, the segmented inclusive graph (SIG) and its HNSW-like variant the hierarchical segmented inclusive graph (HSIG) (see Figure 12). UNIFY assumes a single, ordered attribute. It divides the range of possible attribute values into $s$ segments, s.t., each segment contains approximately the same number of items. The high-level idea of a SIG is to build a proximity graph (Wang et al., 2021b) for each of the $2^{s}-1$ possible combinations of segments, and then merge these graphs into a single graph that contains the $2^{s}-1$ smaller graphs as subgraph. Due to a clever construction algorithm, the SIG can be built without explicitly constructing all $2^{s}-1$ smaller graphs. The SIG is stored as a segmented adjacency list, where outgoing edges are grouped by their destination segment. When querying the SIG, we only traverse the subgraph that correspond to those segments that intersect with the query range. The HSIG is a multi-layered variant of the SIG which is traversed from top to bottom in order to speed-up the query execution. To handle queries with low selectivity, UNIFY supports a pre-filtering variant and for queries with high selectivity, it supports unfiltered ANNS with post-filtering.
Figure 12. 3.3.18) UNIFY’s HSIG index visualized.
3.3.19. UNG
UNG (Cai et al., 2024) (see Figure 13) is a graph-based FANNS index that supports both EMIS and MEMIS filters. It first identifies all distinct label sets in the database and constructs a Label Navigating Graph (LNG), which is a Directed Acyclic Graph (DAC) with one vertex for each distinct label set and a directed edge from label set $a$ to label set $b$ if $a$ is a subset of $b$. The Unified Navigating Graph (UNG) index is built by constructing a proximity graph for each vertex in the LNG, where the vertices of each proximity graph represent the items with the corresponding label set. To unify these proximity graphs, cross-edges are added between vertices in the proximity graphs of label sets $a$ and $b$ if there is a corresponding edge from $a$ to $b$ in the LNG. When querying the UNG, the standard graph traversal is initiated, starting in all proximity graphs whose label sets are minimal supersets of the query label set.
Figure 13. 3.3.19) Visualization of the UNG index.
4. A New Dataset to Benchmark FANNS
Benchmarking FANNS methods requires suitable datasets. Such datasets should contain vectors from a state-of-the-art embedding model (requirement 1), real-world attributes (requirement 2), a set of queries (vectors and filters) together with the pre-computed ground truth (requirement 3), and be publicly available (requirement 4).
The most common approach to evaluate FANNS methods that we observed in scientific literature is to use an ANNS dataset such as SIFT (Laurent Amsaleg, 2025), GIST (Laurent Amsaleg, 2025), MNIST (LeCun, 2025), GloVe (Pennington et al., 2025), UQ-V (Fu, 2025), Msong (Fu, 2025), Audio (Fu, 2025), Crawl (Fu, 2025), Enron (Fu, 2025), BIGANN (Simhadri et al., 2025), Paper (Fu, 2025), Deep1B (Dmitry Baranchuk, 2025), MSTuring (Simhadri et al., 2025), or YandexT2I (Simhadri et al., 2025) and extend it with synthetic, randomly generated attributes (Xu et al., 2020; Wang et al., 2021d; Wang et al., 2024; Yang et al., 2020c; Zhao et al., 2022; Mohoney et al., 2023; Gollapudi et al., 2023; Gupta et al., 2023; Engels et al., 2024; Zuo et al., 2024; Liang et al., 2024; Cai et al., 2024) (failing to meet requirement 2). Another common, though more labor-intensive, approach is to repurpose a dataset not originally intended for FANNS such as TripClick (Rekabsaz et al., 2025), LAION (Schuhmann, 2025), RedCaps (Desai et al., 2025), YouTube-Rgb (Abu-El-Haija et al., 2025), Youtube-Audio (Abu-El-Haija et al., 2025), Words (Farrall, 2025), MTG (Trevor, 2025), or WIT-Image (Datasets, 2025) by manually creating embedding vectors and queries, or by crawling the web to obtain suitable real-world attributes (Patel et al., 2024; Engels et al., 2024; Zuo et al., 2024; Xu et al., 2024; Liang et al., 2024; Cai et al., 2024) (failing to meet requirement 3). Finally, some papers originating from industry rely on proprietary in-house datasets that are not publicly available (Wei et al., 2020; Yang et al., 2020c; Mohoney et al., 2023) (violating requirement 4).
Figure 14. 4) Distribution of the unordered attributes (blue), ordered attributes (green), and set attributes (yellow) in the arxiv-for-fanns-large dataset. We show inverse histograms for authors and submitter since they have many possible values and only few items with a given value. Outliers are omitted for clarity. The has_comment attribute is true for 74.216% of items.
The only two datasets that appear to satisfy all four requirements are MNIST (LeCun, 2025) and Paper (Fu, 2025). The MNIST dataset, where the labels correspond to the digits 0 to 999, and the Paper dataset, where each attribute has at most three possible values, cover only a narrow subset of potential FANNS use cases. Moreover, we observe that datasets with text embeddings from state-of-the-art, transformer-based models such as NV-Embed (Lee et al., 2024; NVIDIA, 2025), LENS (Lei et al., 2025; Lei, 2025), GTE (Li et al., 2023; Alibaba-NLP, 2025), Stella (Zhang et al., 2024; NovaSearch, 2025), and BGE (Chen et al., 2024; of Artificial Intelligence, 2025) are absent from the current ANNS and FANNS benchmarks (not in line with requirement 1). These models typically generate normalized embedding vectors with around 4096 dimensions, representing a fundamentally different FANNS workload compared to existing datasets, which usually contain embeddings with fewer than 1000 dimensions. We therefore argue that an additional datasets with transformer-based embeddings would be highly valuable to the community.
Table 4. 4) Attributes in the novel arXiv dataset.
Name Attribute Type Data type
submitter unordered attribute string
has_comment unordered attribute boolean
number_of_main_categories ordered attribute integer
main_categories set attribute list of strings
number_of_sub_categories ordered attribute integer
sub_categories set attribute list of strings
license unordered attribute string
number_of_versions ordered attribute integer
update_date ordered attribute integer
number_of_authors ordered attribute integer
authors set attribute list of strings
To address this gap, we introduce the arxiv-for-fanns dataset, where each item corresponds to a research paper. The dataset is available in three scales: small (1k items), medium (100k items), and large (over 2.7M items). Each item includes a 4096-dimensional, normalized text embedding of the paper abstract, computed using the stella_en_400M_v5 model (NovaSearch, 2025; Zhang et al., 2024), based on abstracts from the arXiv dataset (Cornell University et al., 2025), thus satisfying requirement 1. Among the top 10 models on the English MTEB leaderboard (Muennighoff et al., 2022), Stella is the most lightweight [^2] As of March 11, 2025 (Benchmark, 2025). Each item also contains 11 real-world attributes (see Table 4) covering all three attribute types (see Section 2.3), with diverse value distributions shown in Figure 14, fulfilling requirement 2. To satisfy requirement 3, each dataset variant includes three sets of 10k queries for EM, R, and EMIS filters (see Section 2.4), along with their corresponding precomputed ground truth. Query vectors are generated by embedding search terms synthesized by GPT-4o (Hurst et al., 2024), while filter values are sampled from the dataset attributes. Figure 15 illustrates the selectivity distribution (as defined in Section 3.1) across queries in arxiv-for-fanns-large. To support unfiltered ANNS evaluation, we also provide ground truth results without any filtering. All three versions of the dataset are publicly available on Hugging Face [^3] https://huggingface.co/datasets/SPCL/arxiv-for-fanns-large , thereby fulfilling requirement 4.
Figure 15. 4) Selectivity of the three query sets.
5. Benchmarking FANNS
We benchmark a selection of FANNS methods on our newly introduced arxiv-for-fanns dataset (see Section 4). Experiments with the medium-scale dataset are conducted on a laptop, while those on the large-scale dataset are executed on a compute cluster.
Table 5. 5.2) Parameters used for benchmarking (found through parameter search, see Section 5.2).
Method Experiment Parameters
ACORN (Patel et al., 2024) all $efs\in\{10,15,20,25,30,50,100,250,500,750\}$
medium, EM filter $M=16,M_{\beta}=24,\gamma=10$
medium, R filter $M=32,M_{\beta}=24,\gamma=12$
medium, EMIS filter $M=16,M_{\beta}=24,\gamma=15$
large, EM filter $M=32,M_{\beta}=32,\gamma=10$
large, R filter $M=32,M_{\beta}=16,\gamma=12$
large, EMIS filter $M=48,M_{\beta}=48,\gamma=15$
CAPS (kmeans) (Gupta et al., 2023) all $m\in\{200,400,1k,5k,10k,20k,40k,60k\}$
medium, EM filter $B=128$
large, EM filter $B=512$
FDANN (stitched) (Gollapudi et al., 2023) all $L\in\{10,20,30,50,100,150,200,300,500,1k\}$
medium, EM filter $R_{\text{small}}=32,L_{\text{small}}=80,R_{\text{stitched}}=48,\alpha=1.1$
medium, EMIS filter $R_{\text{small}}=32,L_{\text{small}}=100,R_{\text{stitched}}=48,\alpha=1.2$
large, EM filter $R_{\text{small}}=96,L_{\text{small}}=100,R_{\text{stitched}}=64,\alpha=1.4$
large, EMIS filter $R_{\text{small}}=64,L_{\text{small}}=80,R_{\text{stitched}}=64,\alpha=1.2$
NHQ (kgraph) (Wang et al., 2024) all $L_{\text{search}}\in\{10,25,50,75,100,150,200,300,400\},\text{iter}=12,M=1.0$
medium, EM filter $K=100,L=80,S=15,R=200,\text{RANGE}=70,PL=200,B=0.5,\text{weight\_search}=1M$
large, EM filter $K=80,L=60,S=10,R=200,\text{RANGE}=60,PL=300,B=0.6,\text{weight\_search}=1M$
NHQ (nsw) (Wang et al., 2024) all $\text{efSearch}\in\{50,100,150,200,300,500,1k,2k,4k\}$
medium, EM filter $M=40,\text{MaxM0}=40,\text{efConstruction}=300,\text{weight\_search}=1M$
large, EM filter $M=40,\text{MaxM0}=60,\text{efConstruction}=150,\text{weight\_search}=1M$
UNG (Cai et al., 2024) all $L_{\text{search}}\in\{10,20,30,40,50,100,150,200,300,500,1k\}$
medium, EM filter $\delta=2,R=24,L=80,\alpha=1.4,\sigma=16$
medium, EMIS filter $\delta=8,R=32,L=100,\alpha=1.4,\sigma=16$
large, EM filter $\delta=6,R=48,L=150,\alpha=1.6,\sigma=24$
large, EMIS filter $\delta=6,R=48,L=150,\alpha=1.4,\sigma=24$
[…]
5.1. Collected Metrics
We focus on the well-established recall vs. queries per second (QPS) plots, which illustrate the trade-off between accuracy and query throughput achieved by each method. We define recall@k as:
[…]
where $\text{knn}_{\text{alg}}$ denotes the set of $k$ nearest neighbors returned by the algorithm, and $\text{knn}_{\text{gt}}$ is the ground truth. In addition, we report the index construction time, peak memory usage during both index construction and query execution, and the index size.
5.2. Parameter Search
Since all FANNS methods considered involve tunable parameters that influence index construction, we perform a dedicated parameter search for each combination of method, dataset scale (medium vs. large), and filter type prior to benchmarking. Given that some methods expose up to eight parameters, a full grid search is computationally infeasible. Instead, we adopt a greedy parameter search strategy, outlined in Algorithm 1. The algorithm takes as input a list of tunable parameters, their corresponding candidate values, and the index of each parameter’s default value inside the list of candidate values. Candidate and default values are selected based on the literature and publicly available implementations of each method. The get_reward() function performs two iterations of building the index with a given parameter configuration and querying it using 505050 randomly sampled queries from the dataset. It computes the recall vs. QPS curve averaged over both iterations and returns the highest QPS achieved at a recall of at least 0.95. If a method does not reach a recall of 0.95, it returns the highest achieved recall instead. The resulting parameter are listed in Table 5.
5.3. Benchmarking Methodology
Once the parameter search is complete, we benchmark each method five times (without warm-up) and report the mean and standard deviation for all measured metrics following scientific benchmarking practice (Hoefler and Belli, 2015). Index construction is performed using all available hardware threads on the respective system (details provided in the next paragraph), while query execution is restricted to a single thread [^4] Our script reports multiple threads during query execution of CAPS and SeRF, however, CPU utilization confirms that only one thread is active and the others are idle., following common practice (Cai et al., 2024; Gupta et al., 2023; Wang et al., 2024). We search for $k=10$ nearest neighbors and report recall@10, which is a standard choice in the literature (Wang et al., 2024; Yang et al., 2020c; Gollapudi et al., 2023; Patel et al., 2024; Cai et al., 2024). All benchmarks use Euclidean distance, as it is supported by all methods under evaluation. Since our transformer-based embeddings are normalized, KNN under Euclidean distance is equivalent to KNN under cosine distance. We exclude any preprocessing, such as sorting items by attribute value or transforming vectors into alternate formats when reporting index construction time.
Algorithm 1 5.2) Greedy Parameter Search
1:params, value_lists, default_indices
2:indices = default_indices
3:P = {p : value_lists[p][indices[p]] for p in params}
4:best_reward = get_reward(P)
5:do_repeat = true
6:while do_repeat do
7:  do_repeat = false
[…]
10:     P = {p : value_lists[p][indices[p]] for p in params}
11:     P[par] = value_lists[par][indices[par] + change]
12:     reward = get_reward(P)
13:     if reward ¿ best_reward then
14:      if reward ¿ (best_reward * 1.01) then
15:        do_repeat = true
16:      end if
17:      best_reward = reward
18:      indices[par] = indices[par] + change
19:     end if
20:   end for
21:  end for
22:end while
23:return {p : value_lists[p][best_indices[p]] for p in params}
5.4. Software and Hardware Configuration
To ensure reproducibility, we encapsulate our benchmarking infrastructure in a Docker container based on Ubuntu 20.04.6, which includes all necessary dependencies for the evaluated methods. Inside the container, we use Python 3.8.10, gcc 9.4.0, and g++ 9.4.0. Benchmarks on the arxiv-for-fanns-medium dataset are conducted on a laptop equipped with 16 GB of RAM and an Intel Core i7-1165G7 CPU with 4 physical cores and 8 hardware threads, running Arch Linux with kernel version 6.15.5-arch1-1. For the arxiv-for-fanns-large dataset, we execute benchmarks on a single node of a compute cluster with 384 GB of RAM and two Intel Xeon Gold 6154 CPUs, each with 18 physical cores and 36 hardware threads, running CentOS Linux 8.
5.5. Remarks on Algorithms
NHQ-kgraph
We exclude the call to optimize_graph() from query execution timing, as it is independent of the number of queries and its cost is amortized when processing large batches.
ACORN
Algorithmically, ACORN only needs to check the filter condition for visited vertices during graph traversal. However, the implementation we benchmark performs this check for all items prior to traversal. Since prior work is inconsistent on whether this cost is included (Cai et al., 2024) or not (Patel et al., 2024), we report two recall vs. QPS curves for ACORN: one including the filtering overhead (labeled “upper bound”) and one excluding it (labeled “lower bound”).
SeRF
Unlike all other methods, SeRF performs index construction and query execution within a single process, without persisting the index to disk. Therefore, we report only the overall peak memory usage, and we omit the index size from our results.
FDANN
The code provided in the FDANN repository (Microsoft, 2025a) stores two versions of the index, each approximately the size of the original dataset. We optimize the code to store only one index at a time, reducing the overall index size to half.
5.6. Results on arxiv-for-fanns-medium
Figure 16. 5.6) Recall@10 vs. QPS plots for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-medium dataset.
Figure 17. 5.6) Index construction time (left), peak memory usage (middle), and index size (right) for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-medium dataset.
Recall vs. QPS (Figure 16)
We observe that ACORN, the only evaluated method applicable to all three filter types, is generally outperformed by more specialized methods. For the EM filter, NHQ (kgraph), FDANN, and UNG achieve the best performance and perform similarly, whereas CAPS, NHQ (nsw), and ACORN are slower at comparable recall levels. The relative performance of ACORN, NHQ (kgraph), and UNG is consistent with previous benchmarks (Cai et al., 2024) on lower-dimensional datasets. FDANN, which showed inferior performance in those earlier studies, excels with our transformer-based embeddings, while NHQ (nsw) and CAPS appear to underperform. For the EMIS filter, previous benchmarks on lower-dimensional datasets (Cai et al., 2024) indicate that UNG outperforms FDANN and ACORN; however, on our transformer-based embeddings, FDANN demonstrates superior performance.
Index Construction Time (Figure 17, left)
We observe that index construction time varies by up to an order of magnitude across methods. CAPS is the fastest to build, followed by FDANN and NHQ (kgraph). SeRF, UNG, and NHQ (nsw) are significantly slower, with ACORN being the slowest across all filters.
Peak Memory Usage (Figure 17, middle)
Except for NHQ (kgraph) and SeRF, which show noticeably higher peak memory usage, differences in memory consumption during index construction and query execution are moderate across methods.
Index Size (Figure 17, right)
The index built by most methods is approximately 1.6 GB, corresponding to the size of the dataset. CAPS and NHQ (kgraph) produce very small indexes, only a few megabytes; however, they need to access the original database vectors during query execution. If we account for the size of the database vectors, the effective index size for these methods is also around 1.6 GB.
5.7. Challenges In Scaling Up the Dataset Size
[…]
Due to this large scale, 444 out of the 777 methods we benchmarked required modifications to their source code to run with the large dataset. UNG (Cai et al., 2024), CAPS (Gupta et al., 2023), and NHQ (KGraph) (Wang et al., 2024) all suffered from int32 overflows during memory allocation. Additionally, CAPS originally allocated a data structure on the stack, which we moved to the heap to support the large dataset. FDANN (Gollapudi et al., 2023), as mentioned in Section 5.5, initially stored two versions of the index, exceeding the local memory of our cluster’s compute node. Addressing the issues enabled benchmarking all 7 algorithms.
5.8. Results on arxiv-for-fanns-large
Figure 18. 5.8) Recall@10 vs. QPS plots for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-large dataset.
Figure 19. 5.8) Index construction time (left), peak memory usage (middle), and index size (right) for the three filter types (EM, R, EMIS) on the arxiv-for-fanns-large dataset.
Recall vs. QPS (Figure 18)
[…]
Index Construction Time (Figure 19, left)
Index construction on the 27×27$\times$27 × larger dataset increased the runtime of most methods by roughly an order of magnitude. Since the large dataset was processed with 9×9$\times$9 × more threads, this suggests that index construction time grows faster than linearly with dataset size or that speedup from additional threads is sublinear. CAPS remains the fastest method, while ACORN continues to lag behind most others. UNG has by far the longest construction time, which may be due to UNG failing to reach the target recall of 0.950.950.95, causing the parameter search to shift toward more expensive configurations.
Peak Memory Usage (Figure 19, middle)
Memory usage of all methods scales roughly linearly with dataset size. Relative differences during index construction and query execution remain similar to those on the medium-scale dataset.
Index Size (Figure 19, right)
UNG stands out by compressing the index to about one fourth of the dataset size without requiring access to the original vectors at query time. As in the medium-scale setting, CAPS and NHQ (kgraph) produce compact indexes of just a few megabytes but require access to the original vectors at query time. For the remaining methods, the index size is approximately equal to the dataset size.
6. Related Work
In the first part of our work, we survey the current state of the FANNS field. While we are not aware of any existing surveys specifically focused on FANNS, several works review the related ANNS domain, such as the surveys by Echihabi et al. (Echihabi et al., 2021) and Han et al. (Han et al., 2023). Other surveys explore prominent application areas of ANNS and FANNS, most notably in the context of retrieval-augmented generation (RAG) for LLMs (Gao et al., 2023) and reasoning language models (RLMs) with integrated RAG capabilities (Besta et al., 2025).
The second part of our paper introduces a novel dataset featuring transformer-based embedding vectors, designed to reflect emerging RAG workloads. This type of data is not represented in the existing landscape of ANNS and FANNS datasets (Laurent Amsaleg, 2025; Pennington et al., 2025; Dmitry Baranchuk, 2025; Simhadri et al., 2025). Our embeddings are generated using the stella_en_400M_v5 model (NovaSearch, 2025; Zhang et al., 2024), a widely used embedding model in LLM research (Besta et al., 2024; Kim et al., 2025). In contrast to our arxiv-for-fanns dataset, end-to-end RAG benchmarks such as the Massive Text Embedding Benchmark (Muennighoff et al., 2022) contain text passages rather than embedding vectors, making them unsuitable for direct evaluation of FANNS methods.
Finally, we benchmark a range of existing FANNS methods (Patel et al., 2024; Gupta et al., 2023; Wang et al., 2024; Cai et al., 2024; Gollapudi et al., 2023; Zuo et al., 2024) on our dataset. While each of these methods includes evaluation results in its respective publication, there is no comprehensive benchmark of FANNS methods on modern, transformer-based embeddings. In the broader ANNS field, several large-scale benchmarks (Aumüller et al., 2020; Aumueller et al., 2025; Wang et al., 2021c) have been established, providing valuable comparisons across methods and datasets.
7. Conclusion
The rapid progress of embedding models for various modalities has intensified the need for fast and accurate FANNS methods. In response, many approaches supporting EM, R, and EMIS filtering have emerged recently. To structure this evolving field, we present a taxonomy and survey of modern FANNS methods. We identify a key gap: the lack of open datasets with transformer-based embeddings enriched with real-world attributes. To address this, we release the arxiv-for-fanns dataset with over 2.7 million 4096-dimensional embedding vectors and 11 real-world attributes. We benchmark several FANNS methods on this dataset and analyze their performance under this challenging workload. No single method excels universally. ACORN is robust and flexible but often outperformed by specialized methods. SeRF performs well for range filters on ordered attributes but lacks categorical filtering support. Filtered-DiskANN and UNG perform well on medium-scale data but fail to scale, highlighting the difficulty of handling high-dimensional transformer-based embeddings. We conclude that the large dimensionality of transformer-based vectors poses a major challenge for FANNS methods. Moreover, we find that efficient parameter tuning remains one of the most pressing open problems in this area.
8. Acknowledgements
We thank the Swiss National Supercomputing Centre (CSCS) for access to their Ault system, which we used to execute our benchmarks. Furthermore, we thank the PLGrid Consortium for access to their Athena system, which we used to create the arxiv-for-fanns dataset. This work was supported by the ETH Future Computing Laboratory (EFCL), financed by a donation from Huawei Technologies. It also received funding from the European Research Council [Uncaptioned image] (Project PSAP, No. 101002047) and from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION).